Engineers continue to work on the equipment…

Tue May 6 14:21:45 PDT 2003 — Engineers continue to work on the equipment problem for the piece of gear in the 9811 hunt group. They have swapped out the “brains” of the unit, and are currently configuring it. The problem should be alleviated soon. -Scott, Support, and OPS (including Russ, Matt K., and John H. in SF)

UPDATE: Repairs have been completed. The 9811 hunt group has been restored to service and is taking connections again. -OPS, Support

Leave a Reply

Your email address will not be published. Required fields are marked *